# Weakly supervised image segmentation for defect-based grading of fresh produce

### v1

@@ Appendix A Implementation details
Table S.2 shows the hyperparamters used for training the Maskformer models. We evaluated the model every five epochs on the validation set and saved the model with the highest Panoptic Quality. All trainings were done on a single NVIDIA RTX A5000 GPU with 24GB memory.
@@
Appendix B Results using four defect categories
@@
Defects Defect masks Old Bruise New Bruise Old Scar New Scar Overall Model PP train val AP IoU AP IoU AP IoU AP IoU mAP mIoU PQ Maskformer Anno. Anno. .$027\pm.060$$011\pm.024$$058\pm.024$$186\pm.107$$039\pm.035$$229\pm.048$$063\pm.026$$376\pm.077$$047\pm.020$$493\pm.018$$471\pm.013$ Maskformer Anno. Anno. .$009\pm.020$$051\pm.061$$063\pm.013$$115\pm.067$$030\pm.021$$199\pm.062$$034\pm.013$$123\pm.061$$034\pm.009$$447\pm.028$$442\pm.021$ Maskformer SAM Anno. .$0.000\pm.0$000 .$0.000\pm.0$000 .$058\pm.031$$166\pm.078$$031\pm.019$$243\pm.038$$064\pm.035$$351\pm.067$$038\pm.014$$488\pm.012$$462\pm.008$ Maskformer SAM Anno. .$0.000\pm.0$000 .$003\pm.006$$071\pm.024$$079\pm.046$$024\pm.018$$178\pm.055$$4.0$014 .$122\pm.035$$034\pm.010$$433\pm.014$$440\pm.007$ Maskformer SAM SAM .$0.000\pm.0$000 .$0.000\pm.0$000 .$061\pm.034$$173\pm.081$$035\pm.020$$247\pm.042$$069\pm.036$$359\pm.067$$041\pm.014$$490\pm.013$$467\pm.007$ Maskformer SAM SAM .$0.000\pm.0$000 .000±.001plus-or-minus.000.001.000$\pm$.$0$001 .$077\pm.026$$083\pm.049$$029\pm.024$$180\pm.058$$043\pm.014$$125\pm.036$$037\pm.010$$434\pm.016$$442\pm.006$ Maskformer SAM2 Anno. .$0.000\pm.0$000 .$0.000\pm.0$000 .$047\pm.019$$144\pm.082$$039\pm.038$$235\pm.086$$062\pm.014$$394\pm.052$$037\pm.007$$490\pm.021$$467\pm.009$ Maskformer SAM2 Anno. .$0.000\pm.0$000 .$021\pm.031$$064\pm.033$$127\pm.083$$027\pm.019$$168\pm.064$$033\pm.012$$094\pm.038$$031\pm.007$$437\pm.017$$437\pm.017$ Maskformer SAM2 SAM2 .$0.000\pm.0$000 .$0.000\pm.0$000 .$049\pm.019$$152\pm.086$$055\pm.065$$242\pm.090$$067\pm.012$$405\pm.053$$043\pm.014$$494\pm.022$$472\pm.011$ Maskformer SAM2 SAM2 .$0.000\pm.0$000 .$020\pm.030$$067\pm.036$$133\pm.086$$038\pm.034$$170\pm.066$$036\pm.013$$098\pm.039$$035\pm.003$$439\pm.018$$442\pm.019$
Figure S.1: Example visualizations of annotated vs. predicted masks using four defect categories. Left: Input Image, Mid: Annotation, Right: Maskformer Predictions (SAM2). Segments are color-coded as follows: Foreground Banana, Background Banana, Old Bruise, Old Scar, New Bruise, New Scar.

### v2

@@ Appendix B Implementation details
Table S.2 shows the hyperparameters used to train the Maskformer models. We evaluated the model every five epochs on the validation set and saved the model with the highest Panoptic Quality. All models were trained on a single NVIDIA RTX A5000 GPU with 24GB of memory.
@@
Appendix C Additional SAM evaluations
Figure S.2: Distribution of IoU values when comparing annotated and SAM-generated defect masks. While the ViT-B and ViT-L variants show a similar distribution, using the largest model size (ViT-H) leads to significantly higher overlap with hand-annotated masks.
Appendix D Results using four defect categories
In this section, we describe our results when using four different defect classes instead of a single joint category. Our results clearly show that while the detection and segmentation of bruises and scars work well, a reliable categorization into one of the four predefined classes is not possible with the current approach and dataset. We hypothesize that this limitation stems from one or more of the following factors:
  1. 1.
    Ambiguous annotation: The classification of defects across the four predefined categories may be subjective to a large degree. The distinction between “old” and ”new“ defects is not always clear-cut, as the transition is gradual. Additionally, distinguishing bruises from scars can be ambiguous, especially for non-experts (see Figure S.1).
  2. 2.
  3. 3.
    The four defect types are unevenly represented in our dataset (37/182/387/834). A more balanced distribution of categories and a larger number of defect samples are likely to improve categorization accuracy. We recommend collecting more examples from the two underrepresented classes for future work.
@@ Appendix D Results using four defect categories
Defects Defect masks Old Bruise New Bruise Old Scar New Scar Overall Model PP train val AP IoU AP IoU AP IoU AP IoU mAP mIoU PQ Maskformer Anno. Anno. .$031\pm.059$$028\pm.043$$043\pm.024$$157\pm.037$$036\pm.018$$225\pm.097$$050\pm.018$$316\pm.054$$040\pm.019$$482\pm.024$$471\pm.023$ Maskformer Anno. Anno. .$030\pm.059$$013\pm.018$$088\pm.039$$114\pm.035$$034\pm.016$$223\pm.083$$017\pm.009$$066\pm.021$$042\pm.022$$436\pm.018$$431\pm.015$ Maskformer SAM-L Anno. .$039\pm.055$$060\pm.080$$034\pm.016$$175\pm.074$$045\pm.034$$179\pm.022$$066\pm.018$$345\pm.046$$046\pm.012$$487\pm.018$$477\pm.013$ Maskformer SAM-L Anno. .$032\pm.049$$031\pm.038$$074\pm.017$$124\pm.051$$038\pm.021$$161\pm.009$$032\pm.014$$104\pm.039$$044\pm.015$$437\pm.018$$435\pm.019$ Maskformer SAM-L SAM-L .$037\pm.053$$063\pm.081$$036\pm.016$$178\pm.075$$045\pm.033$$182\pm.023$$071\pm.019$$353\pm.049$$047\pm.012$$489\pm.018$$478\pm.013$ Maskformer SAM-L SAM-L .$032\pm.049$$027\pm.034$$8.0$018 .$127\pm.051$$036\pm.020$$163\pm.010$$034\pm.015$$108\pm.041$$046\pm.015$$437\pm.018$$436\pm.019$ Maskformer SAM-H Anno. .$019\pm.026$$074\pm.078$$043\pm.020$$157\pm.092$$031\pm.021$$252\pm.063$$056\pm.016$$318\pm.047$$037\pm.014$$495\pm.010$$474\pm.010$ Maskformer SAM-H Anno. .$032\pm.056$$030\pm.029$$093\pm.022$$145\pm.074$$025\pm.029$$230\pm.041$$029\pm.011$100±.050plus-or-minus.100.050.100$\pm$.$0$050 .$045\pm.016$$453\pm.013$$449\pm.009$ Maskformer SAM-H SAM-H .$025\pm.035$$092\pm.105$$049\pm.022$$163\pm.091$$037\pm.024$$264\pm.064$$062\pm.018$$328\pm.047$$043\pm.015$$501\pm.010$$482\pm.011$ Maskformer SAM-H SAM-H .$045\pm.078$$029\pm.027$$106\pm.022$$149\pm.073$$027\pm.028$$239\pm.041$$034\pm.010$$104\pm.051$$053\pm.019$$455\pm.013$$456\pm.010$
Figure S.3: Example visualizations of annotated vs. predicted masks using four defect categories. Left: Input Image, Mid: Annotation, Right: Maskformer Prediction. Segments are color-coded as follows: Foreground Banana, Background Banana, Old Bruise, Old Scar, New Bruise, New Scar.